# EEG-Language Modeling for Pathology Detection

### v2

@@ 3.2.1 Text processing
In order to categorize the textual content in the clinical reports, we employed regular expressions matching for commonly-occurring headings (an overview is provided in Appendix E.2). These enabled the segmentation of individual reports into their respective headings with associated text paragraphs, providing insight into which information in physician reports is encoded in the EEG. We cluster headings into four categories. First, the clinical history cluster contains demographic information in terms of patient age and sex, as well as a brief description of relevant current and/or past pathology and symptoms. The record description cluster includes the physician’s observations of the EEG traces, which describes both normal and abnormal features, often in terms of oscillatory brain activity. The medication cluster contains the patient’s current medication information. Finally, the interpretation cluster summarizes a physician’s thoughts, often including the impression of whether the EEG is normal or pathologically abnormal, as well as a clinical correlation. To investigate whether EEG-language models can learn richer representations by being exposed to a larger variety of text, we also train models by sampling text from all four aforementioned clusters.
@@
Language encoding. Given a sampled paragraph from a clinical report or the LLM-generated summary, we encode this text by relying on the embedding of a special $[cls]$ token which aggregates the representations across all tokens. As such, given a clinical report $\textbf{x}_{l}$, the transformation function $z_{l}$ corresponds to text segmentation or summarization yielding $\tilde{\textbf{x}}_{l}$. Following tokenization, we embed into the $[cls]$ token using $f_{l}$. The resulting text embedding $\textbf{h}_{l}$ may be used for multimodal pretraining.
@@ 3.2.2 EEG processing
EEG preprocessing. EEG data received minimal preprocessing (using MNE [45]). First, the first 10 seconds were removed to reduce the impact of set-up artefacts. Afterwards, a bandpass filter of 0.1-49 Hz was applied and all recordings were resampled to 100 Hz. To reduce the impact of signal artefacts, all EEG signals had their amplitude clipped to $\pm$ 800 $\mu$V. As a large majority of recordings used an average-reference (AR) or linked-ear reference (LE), we only used these recordings and standardized them via transformation to the 20-channel Temporal Central Parasagittal (TCP) montage. To enable fair comparisons between methods, the optimal crop-length out of {5,10,20,30,60} seconds was determined without data-leakage through training and evaluation on subsets of the training data only (Appendix A.2). Based on these results, we used 20 and 60 second crops for EEG-only and EEG-language modeling respectively.
3.2.3 Data subsampling
TUEG contains considerably more abnormal than normal EEGs. As vision-language models have been shown to be sensitive to imbalanced classes [46], we subsample the data to create approximately equal class representation. To do so, we rely on the LLM summaries of reports, which facilitated report classification based on regular expressions due to reoccurring phrasing. If any data of a subject was present in the retrieval or TUAB test set, all their data was excluded from the pretrain subset to avoid data leakage. Further details are provided in Appendix B and resulting sample sizes are shown in Table 1.
Table 1: Dataset sample sizes.
Data subset EEG files Clinical reports Crops (60s) Crops (20s)
Pretrain 15144 11785 270K 813K
TUAB train 2712 Not used 56579 170K
TUAB test 276 Not used 5783 17349
Retrieval test 437 437 8887 26661
Figure 2: EEG-Language models are evaluated on their retrieval ability using top-k accuracy out of 437 patients. Either EEG is retrieved based on a queried clinical report, or vice versa. Error bars indicate standard deviations over five model training runs.
3.2.4 Retrieval
To investigate the information represented in the learned embeddings resulting from EEG-language training, we perform retrieval analyses. Given a medical report describing the patient and their EEG recording, we probe the ability to recover the patient’s EEG by rank-ordering candidate EEG based on embedding similarity. This analysis is also performed in opposite direction, by retrieving the associated report given an EEG recording. As reports refer to an entire recording, EEG embeddings of single crops are averaged. Given that reports generally consist of multiple paragraphs, embeddings of single paragraphs are averaged too. This procedure yields one EEG and report embedding per patient recording, which we use for rank-ordering based on cosine similarity.
The top-K retrieval accuracy, which scores whether the patient’s EEG or report has a rank equal or better than K, is plotted in Figure 2. Many models perform considerably above chance level, indicating the successful generalization of learned multimodal EEG-language information. For both EEG and report retrieval, ELMe,l models tend to outperform ELMl models. However, this discrepancy in performance is particularly prevalent for report retrieval. This is likely due to omission of a text projection head in ELMl, which may therefore lack the flexibility to appropriately separate the EEG reports in latent space. This is likely exacerbated by the current application domain: while we employ a language model finetuned for the medical domain, EEG-relevant text will only represent a small minority of the training data. Therefore, due to the similarity of the clinical reports, they may span only a relatively small part of the latent space. Meanwhile, ELMe,l performs similarly for EEG and report retrieval.
The text sampling also markedly impacts the report retrieval. The clinical history and medication clusters, which contain no direct description of the observed EEG recording, score lowest. While including such information (description cluster) helps considerably, retrieval is particularly effective when a pathology-relevant context is provided (interpretation cluster and LLM summary). This indicates that pathology is a significant source of between-subject variation. Further clear improvements are seen when text from all clusters is sampled, indicating that these clusters contain unique information and that EEG-language modeling can capture multiple dimensions of patient information.
3.2.5 Pathology classification
Next, we study the learned representations in their relevance to clinical pathology. To do so, we first train linear probes to detect pathology on the representations of pretrained models on TUAB under varying amounts of labels (Table 2; Appendix C). Models are trained on single EEG crops, across which we average predictions to obtain a recording-level prediction. We find EEG-language pretraining yields large improvements for pathology detection over EEG-only pretraining, with multimodal models being particularly effective at small sample sizes: at 1% of exposed labels, performance increases reach 7.7% balanced accuracy and 8.4% AUROC. In general, models with the best retrieval performance, also tend to perform better pathology detection. Nevertheless, relative performance differences between EEG-language models are smaller than for retrieval, suggesting that all models obtained representations useful for the narrower task of binary pathology detection.
Language-independent effects of multimodal pretraining. Given the broad outperforming of ELMs compared to EEG-only models, especially for ELMe,l, we further investigate whether the general setup of multimodal pretraining provides inherent benefits. EEG recordings are split into multiple crops, which in turn are all aligned to the same clinical report during pretraining. It follows that EEG crops of a single recording are indirectly aligned to one another to some extent (Figure 1C). We investigated this hypothesis by shuffling reports between patients prior to pretraining. We find that while embeddings of single EEG crops of an untrained encoder are only minimally more similar within-subject than between-subject (ratio of $\sim$1.1x), this effect is much more pronounced after pretraining ELMe,l on correctly paired reports ($\sim$6.3x), and even more so after pretraining on shuffled reports ($\sim$15.7x; figure 3). Linear probing reveals that training ELMe,l on shuffled reports clearly boosts pathology detection over using an untrained encoder and manages to almost match EEG-only pretraining without the need for augmentations (mean accuracies of 73.70%, 81.04%, 83.69%). On the contrary, the ratios for ELMl are close to 1 after training using paired and shuffled reports, with the latter resulting in decreased pathology detection accuracy.
@@ 3.2.5 Pathology classification
Table 2: Pathology detection via linear probing at 1%, 10%, and 100% labeled data of the TUAB training set. The best scores are printed in bold. Standard deviations over five model training runs are included.
@@
Method Text Sampling 1% 10% 100% 1% 10% 100%
BYOL - 72.69$\pm$0.57 79.03$\pm$1.16 79.94$\pm$2.14 78.85$\pm$0.81 86.75$\pm$0.76 88.82$\pm$0.70
VICReg - 71.76$\pm$0.81 79.6$\pm$1.07 82.46$\pm$0.96 78.7$\pm$1.11 86.04$\pm$0.80 88.78$\pm$1.04
ContraWR - 73.30$\pm$1.44 80.72$\pm$1.69 82.44$\pm$1.22 80.30$\pm$1.91 86.67$\pm$1.32 88.44$\pm$1.20
RP - 74.52$\pm$1.06 82.16$\pm$0.38 83.36$\pm$0.42 82.63$\pm$0.87 89.78$\pm$0.43 91.43$\pm$0.34
TS - 74.99$\pm$0.86 82.16$\pm$0.64 84.10$\pm$0.66 82.51$\pm$0.91 89.58$\pm$0.55 91.50$\pm$0.32
CPC - 73.20$\pm$0.79 78.44$\pm$1.00 79.95$\pm$1.49 81.48$\pm$1.02 86.44$\pm$1.07 87.92$\pm$1.14
ELMl Clinical History 76.36$\pm$0.54 79.88$\pm$1.32 82.61$\pm$1.43 84.48$\pm$1.07 87.87$\pm$1.05 89.40$\pm$0.80
ELMl Medication 75.71$\pm$1.14 80.41$\pm$0.77 83.20$\pm$1.17 84.27$\pm$0.92 88.10$\pm$0.86 89.79$\pm$0.76
ELMl Description 79.61$\pm$0.69 81.87$\pm$0.69 83.88$\pm$0.89 87.88$\pm$0.80 89.73$\pm$0.63 90.67$\pm$0.42
ELMl Interpretation 80.57$\pm$0.62 81.98$\pm$1.41 83.08$\pm$1.01 88.86$\pm$0.60 89.82$\pm$0.78 90.53$\pm$0.62
ELMl LLM Summary 82.36$\pm$0.56 83.70$\pm$0.54 84.37$\pm$0.51 90.35$\pm$0.34 90.97$\pm$0.35 91.58$\pm$0.22
ELMl All Clusters 79.07$\pm$0.87 81.07$\pm$0.75 83.18$\pm$0.60 87.12$\pm$0.48 88.61$\pm$0.36 89.78$\pm$0.23
ELMe,l Clinical History 79.86$\pm$0.00 82.71$\pm$0.00 84.13$\pm$0.00 87.61$\pm$0.00 90.72$\pm$0.00 91.81$\pm$0.00
ELMe,l Medication 79.86$\pm$0.00 82.58$\pm$0.00 82.31$\pm$0.00 88.41$\pm$0.00 90.57$\pm$0.00 91.81$\pm$0.00
ELMe,l Description 81.47$\pm$0.29 83.64$\pm$0.54 84.84$\pm$0.91 89.14$\pm$0.53 91.70$\pm$0.19 92.71$\pm$0.14
ELMe,l Interpretation 82.83$\pm$0.35 84.09$\pm$0.52 84.51$\pm$0.58 90.92$\pm$0.35 92.48$\pm$0.31 93.13$\pm$0.27
ELMe,l LLM Summary 82.18$\pm$0.83 83.16$\pm$1.04 83.24$\pm$0.44 90.35$\pm$0.37 91.57$\pm$0.53 92.27$\pm$0.42
ELMe,l All Clusters 82.64$\pm$0.24 84.13$\pm$0.35 85.39$\pm$0.45 90.98$\pm$0.29 92.53$\pm$0.21 93.26$\pm$0.24
@@
Zero-shot pathology detection. Next, we investigate the unique ability of multimodal language modeling to leverage the language modality to perform ‘zero-shot’ classification. Without any explicit labels for downstream training, EEG may be classified by computing its similarity in latent space to text prompts representing the candidate classes. As suggested by [12], we create a prompt ensemble over 21 variations of the phrasing "The EEG is {normal, abnormal}" (Appendix C). Results in table 3 indicate that, despite a small dataset, EEG-language models can reach high levels of zero-shot pathology detection. The best models outperform all linear probes at 1% labels and even match EEG-only models at 100% labels. The clinical history, medication, and description models perform poorly, which follows from these models not being exposed to the explicit phrasing indicating the EEG status as normal or abnormal per se. Their performance can likely be improved by designing appropriate prompts.
Notably, while the ELMe,l models trained on either the interpretation cluster or all clusters both perform well with high consistency, training on LLM summaries resulted in highly variable scores. As the LLM-generated text was considerably more uniform with repetitive phrasing across reports, the lack of variability in combination with limited data may have lead to unstable language representations of the text projector. Meanwhile, we observe the opposite pattern for ELMl, where LLM summaries enabled the only consistently above-chance zero-shot classifier. As with the report retrieval analysis, the fixed text representations of a language model which is not finetuned for EEG is likely inadequate to reliably separate between diverse descriptions of pathological and normal EEG. Meanwhile, the rigid LLM-generated text may have aided in this scenario by consistently yielding divergent text representations with which normal and abnormal EEG may be aligned. It must be noted that the current set-up may have favoured the LLM-generated summaries as they were used to subsample the dataset into a balanced subset. However, in this case it is likely this filtering step also aided other models by omitting some EEGs for which the pathology status was uncertain.
Table 3: Pathology detection via zero-shot classification. The best scores are printed in bold. Standard deviation over five model training runs are included.
Method Text Sampling Balanced Accuracy AUROC (%) F1-Score (%)
ELMl Clinical History 50.00$\pm$0.00 34.34$\pm$1.66 0.00$\pm$0.00
ELMl Medication 50.00$\pm$0.00 74.99$\pm$1.80 0.00$\pm$0.00
ELMl Description 50.00$\pm$0.00 43.70$\pm$1.99 0.00$\pm$0.00
ELMl Interpretation 50.00$\pm$0.00 89.23$\pm$0.31 0.00$\pm$0.00
ELMl LLM Summary 71.98$\pm$0.62 90.92$\pm$0.35 61.77$\pm$1.24
ELMl All Clusters 50.00$\pm$0.00 84.52$\pm$0.57 0.00$\pm$0.00
ELMe,l Clinical History 62.50$\pm$8.74 67.57$\pm$10.41 60.96$\pm$6.13
ELMe,l Medication 51.09$\pm$17.79 52.21$\pm$23.54 50.67$\pm$12.53
ELMe,l Description 64.03$\pm$3.95 71.40$\pm$4.32 63.56$\pm$3.81
ELMe,l Interpretation 82.34$\pm$1.42 91.80$\pm$0.47 80.10$\pm$1.63
ELMe,l LLM Summary 58.87$\pm$15.48 67.98$\pm$20.88 64.32$\pm$9.06
ELMe,l All Clusters 83.16$\pm$1.15 91.91$\pm$0.67 81.25$\pm$1.35

### v3

@@ 3.3.1 Text processing
In order to categorize the textual content in the clinical reports, we employed regular expressions matching for commonly-occurring headings (an overview is provided in Appendix E.3). These enabled the segmentation of individual reports into their respective headings with associated text paragraphs, providing insight into which information in physician reports is encoded in the EEG. We cluster headings into four categories, while filtering out headings which are irrelevant such as information on the EEG system, technical issues, or general disclaimers. First, the clinical history cluster of headings contains demographic information in terms of patient age and sex, as well as a brief description of relevant current and/or past pathology. The record description cluster includes the physician’s observations of the EEG traces, which describes both normal and abnormal features, often in terms of oscillatory brain activity. The medication cluster contains the patient’s current medication information. Finally, the interpretation cluster summarizes a physician’s thoughts, often including the impression of whether the EEG is normal or pathologically abnormal, as well as a clinical correlation. To investigate whether EEG-language models can learn richer representations by being exposed to a larger variety of text, we also train models by sampling text from these four aforementioned clusters.
@@
Language encoding. Given a sampled section from a clinical report or the LLM-generated summary, we encode this text by relying on the embedding of the $[cls]$ token which aggregates the representations across all tokens. As such, given a clinical report $\textbf{x}_{l}$, the transformation function $z_{l}$ corresponds to text segmentation or summarization yielding $\tilde{\textbf{x}}_{l}$. Following tokenization, we embed into the $[cls]$ token using $f_{l}$. The resulting text embedding $\textbf{h}_{l}$ may be used for multimodal pretraining.
@@ 3.3.2 EEG processing
EEG data received minimal preprocessing, with our approach detailed in Appendix C.1. We describe the selection of our pretraining dataset in Appendix C.2, which avoids data leakage by excluding any data of subjects present in any of the evaluation data. The resulting sample sizes are shown in Table 1. To enable fair comparisons between methods, the optimal crop-length for recording-level pathology detection was determined out of {5,10,20,30,60} seconds without data-leakage (Appendix B.2), yielding 20 and 60 second crops for EEG-only and EEG-language modeling respectively. For TUSZ and TUEV evaluations, we additionally pretrain models using 5-second crops, drop subjects which feature in either TUEV or TUSZ-test, and train with lower learning rates (base learning rate of 0.02 for ELM-MIL and 0.1 otherwise). For ELM-MIL we additionally increase the amount of positive EEG crops $N$ to 120.
4 Experimental Results
4.1 Pretraining comparisons
Given the novelty of the studied domain, we extensively investigate the information represented in the learned embeddings resulting from EEG-language training as a function of both text selection and alignment strategy. To this end, we perform retrieval analyses as well as pathology detection using linear probes and zero-shot classification (Figure 2).
@@ 4.1 Pretraining comparisons
Table 2: Pathology detection via zero-shot (ZS) and linear probing at 1%, 10%, and 100% labeled data of the TUAB training set. The (second) best scores are printed (underlined) bold. Standard deviations over five model training runs are included.
@@
Method ZS 1% 10% 100% ZS 1% 10% 100%
SV - 71.36$\pm$1.10 81.06$\pm$0.30 84.13$\pm$0.29 - 79.87$\pm$1.30 89.23$\pm$0.51 91.83$\pm$0.32
BYOL - 72.69$\pm$0.57 79.03$\pm$1.16 79.94$\pm$2.14 - 78.85$\pm$0.81 86.75$\pm$0.76 88.82$\pm$0.70
VICReg - 71.76$\pm$0.81 79.6$\pm$1.07 82.46$\pm$0.96 - 78.7$\pm$1.11 86.04$\pm$0.80 88.78$\pm$1.04
ContraWR - 73.30$\pm$1.44 80.72$\pm$1.69 82.44$\pm$1.22 - 80.30$\pm$1.91 86.67$\pm$1.32 88.44$\pm$1.20
RP - 74.52$\pm$1.06 82.16$\pm$0.38 83.36$\pm$0.42 - 82.63$\pm$0.87 89.78$\pm$0.43 91.43$\pm$0.34
TS - 74.99$\pm$0.86 82.16$\pm$0.64 84.10$\pm$0.66 - 82.51$\pm$0.91 89.58$\pm$0.55 91.50$\pm$0.32
CPC - 73.20$\pm$0.79 78.44$\pm$1.00 79.95$\pm$1.49 - 81.48$\pm$1.02 86.44$\pm$1.07 87.92$\pm$1.14
ELM-MIL $l|e$ 68.86$\pm$7.89 82.53$\pm$1.80 86.38$\pm$0.77 87.62$\pm$0.43 75.23$\pm$9.28 89.88$\pm$1.47 92.92$\pm$0.54 93.52$\pm$0.34
ELM-MIL $e|l$ 79.10$\pm$2.93 83.71$\pm$0.59 84.37$\pm$0.97 85.65$\pm$0.97 87.26$\pm$3.19 92.37$\pm$0.43 93.25$\pm$0.27 93.65$\pm$0.16
ELM-MIL $e,l$ 84.31$\pm$0.57 83.10$\pm$0.56 84.21$\pm$0.82 87.11$\pm$0.76 91.56$\pm$1.31 91.54$\pm$0.44 93.14$\pm$0.24 93.91$\pm$0.17
Retrieval analyses. Given a medical report describing the patient and their EEG recording, we probe the ability to recover the patient’s EEG by rank-ordering candidate EEG based on embedding similarity, as well as vice versa. We average embeddings of EEG and text segments within-modality, yielding one EEG and report embedding per recording, which we use for rank-ordering based on cosine similarity.
Top-K retrieval accuracy, assessing if the correct EEG or report ranks within the top K (Figure 2A,B), shows that many models perform well above chance. This indicates successful generalization of multimodal EEG-language learning. Text sampling significantly impacts report retrieval; clusters lacking direct EEG descriptions (clinical history, medication) score lowest. Including descriptive information improves retrieval, but pathology-relevant contexts (interpretation, LLM summary) are most effective, highlighting pathology as a key source of variation. Combining multiple text clusters further enhances results, suggesting unique information capture and the model’s ability to integrate diverse patient data.
Table 3: Linear probing for abnormality classification on the NMT dataset using 1%, 10%, and 100% labeled training data.
AUROC
Method 1% 10% 100%
BYOL 63.78$\pm$1.70 76.48$\pm$2.10 80.65$\pm$2.50
ContraWR 65.72$\pm$1.01 72.47$\pm$0.95 75.42$\pm$1.01
VICReg 61.57$\pm$1.49 74.19$\pm$0.63 78.50$\pm$1.60
TS 64.90$\pm$0.70 81.36$\pm$1.53 87.08$\pm$1.02
RP 64.92$\pm$0.81 80.42$\pm$1.83 86.50$\pm$2.17
CPC 65.24$\pm$2.06 77.84$\pm$1.12 79.98$\pm$1.60
ELM-MIL 69.49$\pm$2.26 81.42$\pm$1.15 89.77$\pm$0.21
ELMe,l models tend to outperform ELMl models, particularly for report retrieval. This is likely due to omission of a text projection head in ELMl, which may therefore lack the flexibility to appropriately separate the EEG reports in latent space. Due to the benefit of pretraining when sampling from the four text clusters, we pretrain our ELM-MIL models in this manner only. We observe that our MIL extension further improves retrieval performance, benefiting from sampling multiple positives jointly ($e,l$) and performing bidirectional alignment. These results indicate for the additional flexibility of a MIL-based approach to aid in multimodal alignment, supporting the hypothesis that not all EEG and text pairs are equally informative.
Table 4: Crop-level performance on TUAB (80% labels) of supervised and self-supervised methods using different training datasets and EEG encoders. Model sizes refer to trainable parameters for the EEG encoders.
Methods Fine tuned Model Size B. Acc. AUROC
SPaRCNet Y 0.79M 78.96$\pm$0.18 86.76$\pm$0.12
ContraWR Y 1.6M 77.46$\pm$0.41 84.56$\pm$0.74
CNN-Transformer Y 3.2M 77.77$\pm$0.22 84.61$\pm$0.13
FFCL Y 2.4M 78.48$\pm$0.38 85.69$\pm$0.51
ST-Transformer Y 3.5M 79.66$\pm$0.23 87.07$\pm$0.19
BIOT Y 3.2M 79.59$\pm$0.57 88.15$\pm$0.43
CEReBrO Y 3.58M 79.40$\pm$0.19 87.49$\pm$0.33
CEReBrO Y 40.0M 81.29$\pm$0.15 88.67$\pm$0.06
CEReBrO Y 85.2M 81.67$\pm$0.23 89.16$\pm$0.38
LaBraM-Base Y 5.8M 81.40$\pm$0.19 90.22$\pm$0.09
LaBraM-Large Y 46M 82.26$\pm$0.15 91.27$\pm$0.05
LaBraM-Huge Y 369M 82.58$\pm$0.11 91.62$\pm$0.16
ELM-MIL N 0.93M 84.42$\pm$0.21 91.44$\pm$0.11
Pathology detection. Next, we study the learned representations in their relevance to clinical pathology by training linear probes on TUAB (Figure 2C). Here we observe similar patterns for recording-level classification, with bidirectional ELM-MIL scoring best. Interestingly, good classification was possible with many of the models, causing us to investigate whether the strategy of sub-unit multimodal modeling provides inherent benefits. We provide this additional set of analyses in Appendix A.2, which indicates that our sub-unit alignment strategy promotes the encoding of between-subject information even in the absence of semantically relevant text. This allows ELMe,l to nearly match the best EEG-only pretraining strategy for pathology detection when reports are randomly shuffled.
Table 5: Linear probing for seizure and event classification on the TUSZ and TUEV datasets respectively. The (second) best SSL scores are printed (underlined) bold. Standard deviations over runs are included.
AUROC TUSZ TUEV
Method 1% 10% 100% 100%
SV 84.00$\pm$0.90 87.33$\pm$0.93 90.38$\pm$0.39 86.28$\pm$0.78
BYOL 77.45$\pm$3.14 86.60$\pm$0.79 88.89$\pm$1.02 82.40$\pm$1.60
ContraWR 68.86$\pm$3.13 82.75$\pm$1.66 85.68$\pm$0.74 84.11$\pm$1.64
VICReg 69.61$\pm$1.82 82.01$\pm$0.93 86.17$\pm$0.98 83.26$\pm$1.83
TS 78.60$\pm$1.90 87.63$\pm$0.58 89.77$\pm$0.72 84.86$\pm$1.32
RP 64.41$\pm$2.50 76.72$\pm$1.37 79.52$\pm$1.26 78.95$\pm$1.59
CPC 72.01$\pm$1.00 81.77$\pm$2.24 85.91$\pm$2.00 81.94$\pm$1.75
ELM-MIL 78.98$\pm$5.18 88.98$\pm$0.86 91.51$\pm$0.33 87.69$\pm$1.01
@@
Next, we investigate the unique ability of multimodal language modeling to leverage the language modality to perform ‘zero-shot’ classification (Figure 2D). Without any explicit labels for downstream training, EEG may be classified by computing its similarity in latent space to text prompts representing the candidate classes. As suggested by [12], we create a prompt ensemble over 21 variations of the phrasing ”The EEG is normal, abnormal” (Appendix D). Despite a small dataset, EEG-language models can reach high levels of zero-shot pathology detection if: 1) they include a text projector, 2) include at least the clinician interpretation, and 3) do bidirectional alignment. These models, especially ELM-MILe,l, score highly consistent across different weight initializations, suggesting sound alignment between modalities. Pretraining using exclusively one of the other text clusters yielded poor or inconsistent performance, which follows from these models not being exposed to the explicit phrasing indicating the EEG status as normal or abnormal per se. Their capability likely can be improved by designing appropriate prompts. Having identified the requirements for stable alignment, we focus on the ELM-MILe,l model for EEG-only baselines due to its consistent performance across tasks and seeds.
Table 6: Ablation studies (Means over five training runs). Ret: EEG Retrieval (Top-10 accuracy), LP: Linear Probe (Balanced accuracy at 100%), ZS: Zero-shot classification (F1).
(a) Aggregation
Method Ret. LP ZS
Max+InfoNCE 3.9 77.5 43.2
Attn+InfoNCE 8.3 84.9 17.5
Sum+InfoNCE 24.7 86.0 78.8
MIL-InfoNCE 27.1 87.1 82.1
(b) Positive EEG Samples
N Ret. LP ZS
2 19.8 85.9 78.8
4 21.8 85.9 79.2
8 25.3 86.5 80.0
16 26.5 86.9 78.5
32 27.1 87.1 82.1
(c) Positive Text Samples
M Ret. LP ZS
2 28.1 85.8 80.4
4 27.1 86.0 80.4
8 27.1 87.1 82.1
4.2 Baseline comparisons
4.2.1 Abnormality classification
We compare our ELM-MIL models on the TUAB dataset to six EEG-only methods (Table 2). We pretrain using these methods on the same pretraining dataset as ELMs and, together with the supervised baseline, also an identical EEG encoder architecture. This enabled accurate inference on the effectiveness of the pretraining strategy per se. We find EEG-language pretraining yields large improvements for pathology detection over EEG-only pretraining, with multimodal models being particularly effective at small sample sizes: at 1% of exposed labels, performance increases reach 8.7% balanced accuracy and 9.7% AUROC. To validate our results out-of-distribution, we evaluate learned representations on the NMT dataset without finetuning (Table 3). We find ELM-MILe,l (henceforth ’ELM-MIL’) to still perform well, especially with few labels, where they yield an improvement of at least 3.8% AUROC.
Next, we further contextualize our results in the literature. By probing ELM-MIL representations for 10 second crop-level abnormality classification on TUAB with 80% labels we are able to compare to previously reported evaluations [20, 11, 47]. However, we note that methods use significantly different pretraining data and architectures, which complicates interpretation. Nevertheless, compared to supervised and self-supervised methods, ELM-MIL scores highest and improves over LaBraM-Huge +0.83% on average across scores (Table 4). This is despite LaBraM-Huge requiring finetuning, being pretrained on a large number of datasets (including TUEG), and featuring 369M parameters compared to 0.9M of our EEG encoder. This is likely due to the clinical specificity of ELMs, enabling much smaller models that do not need to be finetuned. As a result, our methodology is also significantly cheaper and faster to train (50 epochs in under 12 hours with 24GB of memory).
4.2.2 Seizure and event classification
We additionally investigated classification performance based on single, five second EEG crops. Although recording-level classification was the focus of our work, we find ELM to score better than the investigated EEG-only methods on both TUSZ and TUEV as well (Table 5). Due to the smaller size of TUEV, we evaluate only using 100% labels. We provide per-class performance in Appendix A.1.
4.2.3 Further Ablation Analyses
Whereas for InfoNCE the temperature parameter sets the relative focus across negative samples [48], for MIL-InfoNCE it does so too for positive samples. We therefore test the sensitivity of our methods to the parameter (Figure 3). We find that MIL-InfoNCE is more robust to changes of $\tau$ for pathology detection, while retrieval performance can be further improved by lowering $\tau$. This may be explained as retrieval being subject-based rather than class-based (see Appendix 12). Moreover, performance increases from $\tau<1$ indicates the utility of this additional hyperparameter of InfoNCE, which is absent in NCE.
Figure 3: Model comparisons across EEG retrieval and pathology detection under different values of the temperature parameter $\tau$.
We perform additional ablations to investigate crucial aspects of the ELM-MIL $e,l$ model. First, we find that additional positive EEG and text samples improve downstream performance (Table 6). We additionally ablate the aggregation method for positive samples and find MIL-InfoNCE to outperform considered alternatives. We compare to aligning only the most similar positive sample (denoted Max+InfoNCE), using attention to create a weighted average across positive samples based on similarity values (Attn+InfoNCE; [49]), as well as taking the sum instead of mean across log-probabilities (Sum+InfoNCE). The latter does not account for the varying amount of text and EEG crops across subjects.